OpenStudy (anonymous):

Which of the following is the solution of (logx - 3) 0.001 = -3 a.x = 5 b. x = 7 c. x = 13 d. x = 15

OpenStudy (anonymous):

is 0.001 being multiplied by log(x) - 3?

OpenStudy (anonymous):

this is how the question was written for me log x - 3 0.001 = -3

OpenStudy (luigi0210):

Rewrite it as an exponent (x-3)^-3=.001

OpenStudy (anonymous):

\[\large \log_{x-3}(0.001)=-3\]

OpenStudy (anonymous):

@satellite73 yes

OpenStudy (anonymous):

what @Luigi0210 said

OpenStudy (anonymous):

and what @genius12 is about to say

OpenStudy (anonymous):

@satellite73 lol@u

OpenStudy (anonymous):

@Luigi0210 so after I write as an exponent (x-3)^-3=.001 what do I do next?

OpenStudy (luigi0210):

multiply it by -1/3, both sides

OpenStudy (luigi0210):

* exponent wise

OpenStudy (anonymous):

so I will get (x-3)=.001^-1/3?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Yup, now what's .001^-1/3?

OpenStudy (anonymous):

I typed it into my calculator and got an error @Luigi0210

OpenStudy (anonymous):

I actually reentered it and ended up with 10 but thats not one of the options :/ @Luigi0210

OpenStudy (anonymous):

oh no

OpenStudy (luigi0210):

Yes, but we are not finished :) Now we have: x-3=10

OpenStudy (anonymous):

\(10^{-3}=0.001\) and so \(x-3=10\) making \(x=13\)

OpenStudy (luigi0210):

thank you @satellite73 xD

OpenStudy (anonymous):

O.O

OpenStudy (luigi0210):

Yup, that simple :)

OpenStudy (anonymous):

hahah thank you :) I appreciate the help! @Luigi0210 @satellite73 @genius12

OpenStudy (luigi0210):

Anytime :D